# AV2AV: Direct Audio-Visual Speech to Audio-Visual Speech Translation with Unified Audio-Visual Speech Representation

### v1

@@ Abstract
This paper proposes a novel direct Audio-Visual Speech to Audio-Visual Speech Translation (AV2AV) framework, where the input and output of the system are multimodal (i.e., audio and visual speech). With the proposed AV2AV, two key advantages can be brought: 1) We can perform real-like conversations with individuals worldwide in a virtual meeting by utilizing our own primary languages. In contrast to Speech-to-Speech Translation (A2A), which solely translates between audio modalities, the proposed AV2AV directly translates between audio-visual speech. This capability enhances the dialogue experience by presenting synchronized lip movements along with the translated speech. 2) We can improve the robustness of the spoken language translation system. By employing the complementary information of audio-visual speech, the system can effectively translate spoken language even in the presence of acoustic noise, showcasing robust performance. To mitigate the problem of the absence of a parallel AV2AV translation dataset, we propose to train our spoken language translation system with the audio-only dataset of A2A. This is done by learning unified audio-visual speech representations through self-supervised learning in advance to train the translation system. Moreover, we propose an AV-Renderer that can generate raw audio and video in parallel. It is designed with zero-shot speaker modeling, thus the speaker in source audio-visual speech can be maintained at the target translated audio-visual speech. The effectiveness of AV2AV is evaluated with extensive experiments in a many-to-many language translation setting. The demo page is available on choijeongsoo.github.io/av2av.
@@ 2.1 Spoken Language Translation
Neural Machine Translation (NMT) [1, 2, 3, 4, 5, 6] has achieved maturity in the modality with the richest data, text. Given that audio modality enables the translation between languages that have no writing systems and the more improved dialogue experience, speech-based translation has emerged. Speech-to-Speech Translation (A2A) [7, 42] aims to translate speech from one language to the semantically consistent speech of another language. Early A2A works began with a cascaded approach [57, 58, 59] by sequentially performing ASR, NMT, and TTS. Subsequently, A2T [39, 38, 40, 33] works were proposed to merge the ASR and NMT stages, and allowed a two-stage A2A system. Recently, it has evolved even into a direct approach [60, 8, 43, 61, 11] that can directly translate speech without relying on intermediate representation.
Despite the recent success of audio-based speech translation, multimodal (i.e., audio and visual) speech translation is in its very early stages. One line of research focuses on language translation between audio input and audio-visual output, namely Speech to Audio-Visual Speech Translation (A2AV). They [15, 16, 17, 18, 19] utilized a 4-stage cascaded approach by chaining ASR, NMT, TTS, and TFG. Specifically, audio from the source video is first transcribed into text with an ASR [20]. Then, the text is translated from the source language to the target language through NMT [23, 4]. The translated text is synthesized into speech using TTS [24, 25], and finally, the face is synthesized from the synthesized speech using TFG [27]. Although the cascaded approach can benefit from the advantage of advancement made in each of the subsystems, they might suffer from slow inference time, domain mismatch, error propagation between the subsystems, and high maintenance costs. Most recently, AV-TranSpeech [62] first proposed direct Audio-Visual Speech-to-Speech Translation (AV2A), whose input is now audio-visual speech and output is audio. By using AV inputs, they showed that the robustness of the translation system on acoustic noise can be improved [63, 64, 65, 66]. Also, they tried to solve the insufficient AV2A training data by bringing pre-trained weights for each modal stream from different pre-trained models [52, 43].
@@ 3.3 Zero-shot Audio-Visual Renderer
Length Predictor. As the output audio and visual streams are expected to be synchronized, we only need one common duration modeling for the two output streams. To this end, a length predictor is employed to predict the duration of each AV speech unit. The predicted duration is utilized for synthesizing both audio and visual outputs. Similar to that of TTS [25], it consists of two 1D convolution layers with one classifier and is trained with Mean Squared Error loss measured between the prediction and ground truth duration of each unit, as in [8, 9]. The translated AV speech units are repeated by the predicted duration before passing the vocoder and face renderer.
@@ 4.1 Dataset
For training the vocoder, we use LRS3 [87] for En and mTEDx [89] for other languages. For the face renderer, we follow the previous works [27] and utilize the widely used LRS3 [87] dataset.
@@ 4.2 Evaluation Metrics
We evaluate the translation quality and the generation quality of both audio and video. For the translation quality, we employ the BLEU score [99, 100] by transcribing the transcription using the off-the-shelf ASR models following [11]. To measure the generation quality of video, we adopt metrics used for TFG. This includes Fréchet Inception Distance (FID) [101] to measure visual quality, and LSE-C and LSE-D [27] to measure the audio-visual synchronization. Also, we conduct a Mean Opinion Score test to measure the naturalness of each respective modality, and the realness of the video.
@@ 4.3 Implementation Details
For training the vocoder, we update 1M steps using a batch size of 16 on each language dataset. For training the face renderer, the AV speech units are embedded into 512 dimensional features at 25fps, which are channel-wise concatenated with the corresponding identity features and decoded by the face decoder. It is trained for 35k steps.
Table 2: Translation quality (ASR-BLEU) comparison with baseline systems for X-En and En-X directions. Methods (A1, A2, A3, A4) use one ASR, AVSR, A2T, and AV2T model for X-En, and 4 different models for En-X directional translations. The methods (A5, A6, A7, A8, A9) use a single model for all 8 directional translations. A2A [11] method does not have a vocoder to generate Portuguese speech.
@@
Table 2: Translation quality (ASR-BLEU) comparison with baseline systems for X-En and En-X directions. Methods (A1, A2, A3, A4) use one ASR, AVSR, A2T, and AV2T model for X-En, and 4 different models for En-X directional translations. The methods (A5, A6, A7, A8, A9) use a single model for all 8 directional translations. A2A [11] method does not have a vocoder to generate Portuguese speech.

### v2

@@ Abstract
This paper proposes a novel direct Audio-Visual Speech to Audio-Visual Speech Translation (AV2AV) framework, where the input and output of the system are multimodal (i.e., audio and visual speech). With the proposed AV2AV, two key advantages can be brought: 1) We can perform real-like conversations with individuals worldwide in a virtual meeting by utilizing our own primary languages. In contrast to Speech-to-Speech Translation (A2A), which solely translates between audio modalities, the proposed AV2AV directly translates between audio-visual speech. This capability enhances the dialogue experience by presenting synchronized lip movements along with the translated speech. 2) We can improve the robustness of the spoken language translation system. By employing the complementary information of audio-visual speech, the system can effectively translate spoken language even in the presence of acoustic noise, showcasing robust performance. To mitigate the problem of the absence of a parallel AV2AV translation dataset, we propose to train our spoken language translation system with the audio-only dataset of A2A. This is done by learning unified audio-visual speech representations through self-supervised learning in advance to train the translation system. Moreover, we propose an AV-Renderer that can generate raw audio and video in parallel. It is designed with zero-shot speaker modeling, thus the speaker in source audio-visual speech can be maintained at the target translated audio-visual speech. The effectiveness of AV2AV is evaluated with extensive experiments in a many-to-many language translation setting. Demo page is available on choijeongsoo.github.io/av2av.
Equal contribution. Corresponding author. This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2022R1A2C2005529), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.2022-0-00124, Development of Artificial Intelligence Technology for Self-Improving Competency-Aware Learning Capabilities), and BK21 FOUR (Connected AI Education & Research Program for Industry and Society Innovation, KAIST EE, No. 4120200113769).
@@ 2.1 Spoken Language Translation
Neural Machine Translation (NMT) [1, 2, 3, 4, 5, 6] has achieved maturity in the modality with the richest data, text. Given that audio modality enables the translation between languages that have no writing systems and the more improved dialogue experience, speech-based translation has emerged. Speech-to-Speech Translation (A2A) [7, 42] aims to translate speech from one language to the semantically consistent speech of another language. Early A2A works began with a cascaded approach [57, 58, 59] by sequentially performing ASR, NMT, and TTS. Subsequently, A2T [39, 38, 40, 33] works were proposed to merge the ASR and NMT stages, and allowed a two-stage A2A system. Recently, it has evolved even into a direct approach [60, 8, 43, 61, 11] that can directly translate speech without relying on intermediate text representation.
Despite the recent success of audio-based speech translation, multimodal (i.e., audio and visual) speech translation is in its very early stages. One line of research focuses on language translation between audio input and audio-visual output, namely Speech to Audio-Visual Speech Translation (A2AV). They [15, 16, 17, 18, 19] utilized a 4-stage cascaded approach by chaining ASR, NMT, TTS, and TFG. Specifically, audio from the source video is first transcribed into text with an ASR [20]. Then, the text is translated from the source language to the target language through NMT [23, 4]. The translated text is synthesized into speech using TTS [24, 25], and finally, the talking face video is synthesized from the synthesized speech using TFG [27]. Although the cascaded approach can benefit from the advantage of advancement made in each of the subsystems, they might suffer from slow inference time, domain mismatch, error propagation between the subsystems, and high maintenance costs. Most recently, AV-TranSpeech [62] first proposed direct Audio-Visual Speech-to-Speech Translation (AV2A), whose input is now audio-visual speech and output is audio. By using AV inputs, they showed that the robustness of the translation system on acoustic noise can be improved [63, 64, 65, 66]. Also, they tried to solve the insufficient AV2A training data by bringing pre-trained weights for each modal stream from different pre-trained models [52, 43].
@@ 3.3 Zero-shot Audio-Visual Renderer
Length Predictor. As the output audio and visual streams are expected to be synchronized, we only need one common duration modeling for the two output streams. To this end, a length predictor is employed to predict the duration of each AV speech unit. The predicted duration is utilized for synthesizing both audio and visual outputs. Similar to that of TTS [25], it consists of two 1D convolution layers with one classifier and is trained with Mean Squared Error loss measured between the prediction and ground truth duration of each unit in the logarithmic domain, as in [8, 9]. The translated AV speech units are repeated by the predicted duration before passing the vocoder and face renderer.
@@ 4.1 Dataset
For training the vocoder, we use LRS3 [87] for En and mTEDx [89] for other languages. For the face renderer, we follow [27] and utilize the widely used LRS3 [87] dataset.
@@ 4.2 Evaluation Metrics
We evaluate the translation quality and the generation quality of both audio and video. For the translation quality, we employ the BLEU score [99, 100] by transcribing the audio using the off-the-shelf ASR models, following [11]. To measure the generation quality of video, we adopt metrics used for TFG. This includes Fréchet Inception Distance (FID) [101] to measure visual quality, and LSE-C and LSE-D [27] to measure the audio-visual synchronization. Also, we conduct a Mean Opinion Score test to measure the naturalness of each respective modality, and the realness of the video, which can be found in Supplementary.
@@ 4.3 Implementation Details
For training the vocoder, we update 1M steps using a batch size of 16 on each language dataset. The face renderer is trained for 35k steps with a batch size of 64.
@@
Table 2: Translation quality (ASR-BLEU) comparison with baseline systems for X-En and En-X directions on MuAViC. Methods (A1, A2, A3, A4) use the same model for X-En, and 4 different models for En-X directional translations. The methods (A5, A6, A7, A8) use a single model for all 8 directional translations. A2A [11] method does not have a vocoder to generate Portuguese speech.
@@
4.4 Baseline Models
Since there is no previous method that can perform AV2AV, we compare the proposed method with the recently proposed direct AV2A method, AV-TranSpeech [62] on LRS3-T dataset. Moreover, we compare the translation performance with different cascaded methods on MuAViC [33] dataset. They are built from state-of-the-art off-the-shelf pre-trained models: AVSR [33], ASR [33], AV2T [33], A2T [33], NMT [5], TTS [105], and TFG [27]. Please note the objective of the comparisons with the cascaded method is not to achieve state-of-the-art performance, but rather to assess the extent to which the performance of the proposed system can be attained through the direct strategy.
The AVSR and ASR models [33] are trained on the multilingual audio-visual corpus, covering 1,200 hours of 9 languages. The AV2T and A2T models [33] support X-to-En language translations with one model while multiple unidirectional models are used for En-to-X translations. Therefore, we use one pre-trained AV2T or A2T model for {Es, Fr, It, Pt}-En translation, while we use 4 individual models for En-Es, En-Fr, En-It, and En-Pt translations. The NMT model [5] is trained on the many-to-many text translation corpus comprising 7.5B sentences for 100 languages. The TTS model [105] is a recent state-of-the-art multilingual TTS model by Team Coqui. The TFG model [27] is one of the popular models that can generate arbitrary identities and languages, trained on 233 hours of audio-visual data.
4.5 Experimental Results
4.5.1 Comparisons with State-of-the-art
We compare the AV2A performance with the state-of-the-art direct AV2A method, AV-Transpeech [62], by using different amounts of finetuning data, in Table 1. The results show that the proposed method is much more effective than AV-Transpeech, especially in the low-resource setting. As the proposed method can be pre-trained using A2A dataset to perform with different input modalities, the pre-trained model can be directly applied to AV2A without finetuning. In contrast, since AV-Transpeech tries to bring different pre-trained models (i.e., AV-HuBERT [52] and A2A [43] models) to initialize different parts of their model, the finetuning with the AV2A dataset is mandatory. Even when no finetuning is applied, the proposed method outperforms AV-Transpeech that finetuned on 10hr of AV2A dataset, and achieves comparable performances with the model finetuned on 30hr of dataset. If we finetune the proposed model with the same amount of AV2A dataset as AV-Transpeech, it outperforms AV-Transpeech at all data sizes. This shows the effectiveness of the proposed strategy of learning from A2A dataset by using unified audio-visual speech representations. Fig. 4 shows the translated results of the proposed AV2AV framework. We show both audio and visual outputs where the audio is transcribed into text through ASR.
4.5.2 Comparisons with Cascaded Approaches
Table 4.3 shows the BLEU score of different systems for different language pairs. The analysis of results is as follows.
(A2, A4 vs. A8) By comparing the methods that can perform AV2AV, the proposed method (A8) shows comparable results with the text-based translation systems trained on 7.5B parallel data (A2, A4), even though it is trained without using any text supervision. This demonstrates that AV speech units encompass sufficient linguistic content, allowing the training of a multilingual AV2AV translation model solely using the discretized AV speech units [^3] Moreover, the proposed method has a faster inference time. On average, the proposed AV2AV required 1.36s, whereas the cascaded system (A2) took 2.42s to process 2.03s video (i.e., audio-visual)..
(A1, A2 vs. A3, A4 vs. A5 vs. A6, A8) We can find that by reducing the cascading stages and shortening the processing time, the BLEU scores become slightly worse overall. This result is attributed to the powerful performance of each advanced subsystem, which benefits from a large-scale dataset. However, we can clearly find that the proposed direct multimodal speech-to-speech translation systems are comparable with the cascaded systems with just a single model. Specifically, the proposed method achieves better performance than the 3-stage cascaded systems (A3, A4) in the X-to-En translation direction. Please note that A1, A2, A3, and A4 use one model for X-to-En and 4 different models to perform En-to-X translation, while the proposed method uses just a single trained model to perform all 8 translation directions in the table. Another important thing is that A5 and the proposed systems are textless so that they can be utilized for the languages that have no writing systems [9], while A1, A2, A3, and A4 cannot be applied to such languages as they rely on text modalities.
(A5 vs. A6, A7, A8) By comparing the proposed method with A2A method (A5), the proposed A2AV method (A6) shows better performances in overall. Moreover, We can confirm that even though the proposed method is pre-trained with the same dataset as A5, the proposed method can be utilized for diverse tasks, including A2AV, V2AV, and AV2AV. This is because the proposed method is trained on unified audio-visual speech representations, eliminating the need for additional training to operate with different modalities. In contrast, A5 utilized audio speech representations during training, so it requires additional finetuning to accept different modal inputs.
  Method X-En
Es-En Fr-En It-En Pt-En
AV-HuBERT [52] 17.63 16.83 21.74 22.22
mAV-HuBERT 26.57 31.27 23.24 24.51
Table 4: Ablation study on multilingual translation performance of English-trained AV-HuBERT and multilingual AV-HuBERT.
  ID Method LSE-C $\uparrow$ LSE-D $\downarrow$ FID $\downarrow$
$\bullet$ Ground Truth
C1 GT Audio-Visual 7.61 6.90 -
[…]
C2 GT Audio + TFG 8.14 6.68 3.56
C3 GT Text + TTS + TFG 7.36 7.06 3.56
[…]
C4 GT AV Speech Unit 7.91 6.65 3.18
Table 5: Performance comparisons of each renderer on LRS3.
4.5.3 Analysis of Robustness to Acoustic Noise
In Table 3, we evaluate the noise robustness of different systems. Following [83], we randomly perturb the input speech with the sampled babble noise from the test set of [104] with varying SNR levels (-5, 0, 5, 10dB). By comparing audio-only input systems with multimodal systems (B1, B3, B5 vs. B2, B4, B7), we can confirm the benefits of using AV inputs by obtaining robust performances under acoustic noises. These results show the importance of AV input systems to achieve robust performance in practical situations. The visual input system (B6) is not affected by acoustic noise and even achieves better performance than audio-only systems (B1, B3, B5) under a severely noisy environment (i.e., -5 dB). As one trained AV2AV model can be employed for diverse tasks, we can choose the input modalities appropriate to a given situation.
4.5.4 Effectiveness of the mAV-HuBERT
In order to evaluate the effectiveness of our multilingual AV-HuBERT (mAV-HuBERT) compared to the original English-trained AV-HuBERT of [52], we compare the BLEU scores on MuAViC [33] X-En translation directions. Table 4 shows the comparison results. The results clearly show that we can improve the multilingual spoken language translation performance largely by pre-training the model with a large-scale multilingual AV dataset (i.e., 7,011 hours with over 100 languages). Conversely, AV-HuBERT trained exclusively in English struggles to accurately capture multilingual speech information, resulting in lower translation performance across all translation directions.
4.5.5 Quantitative Evaluation of AV-Renderer
In Table 4.5.2, we evaluate the visual quality (FID), and degree of synchronization between the generated audio and visual streams (LSE-C and LSE-D). To focus on the performance of renderers, we utilize ground truth audio, text, and AV speech units to generate audio or video for each system. Comparing (C1, C2, C3, C4), the results indicate that the proposed AV-Renderer can generate well-synchronized audio-visual outputs by synthesizing them in parallel, with shared duration information. In contrast, C3 loses accurate synchronization by cascading TTS and TFG, where system error could be propagated between generated samples.
In Table 6, we evaluate the effectiveness of incorporating zero-shot speaker modeling into the vocoder. To this end, we first compare the BLEU and voice similarity (SIM) using WavLM-TDNN [72, 106] between a single speaker-trained vocoder [11] and our zero-shot speaker vocoder. The results confirm that by employing speaker embedding, we do not lose the translation performance, and we can improve the voice similarity (SIM). Please note that this is a key component for seamless speech-to-speech translation. To understand the SIM score, we also report that of a popular multi-speaker TTS, YourTTS [46], and we can confirm that the proposed zero-shot speaker vocoder has sufficient ability to maintain the speaker’s voice.
  Vocoder BLEU SIM
YourTTS [46] - 0.266
[…]
Zero-Shot Speaker 26.57 0.222
Table 6: Ablation study using different vocoders on MuAViC Es-En data. To check the performance of our vocoder, we also report the SIM of a popular multi-speaker TTS, YourTTS [46].
5 Conclusion
In this paper, we proposed a novel direct Audio-Visual Speech to Audio-Visual Speech Translation (AV2AV) framework. The proposed AV2AV can translate spoken languages in a many-to-many setting without text. By employing multimodal inputs, we can improve the robustness of the system to the acoustic noises. By using multimodal outputs, we can improve the dialogue experience performed in virtual scenarios with no discomfort. The effectiveness of the proposed AV2AV is evaluated with extensive experiments.